Image: Scenes from around campus

Past System & Network Status

Date Issue Time of Outage Time restored Scheduled Resolution
Year 2010
         
7/22 Some DAC emplysees’ network access affected because IT Department replaced a power supply on a switch 7:00AM 7:05AM Yes Power supply replaced
 7/02

The PROMT system is temporarily unavailable due to faulty patch

10:00AM 11:10PM Removed a faulty patch
7/01 The PROMT system is temporarily unavailable due to malfunction of an application server 8:20AM 8:40AM   Rebooted Server
 6/18

Installation of Emergency Generator – Power shut down at District

7:00AM 7:30PM  Yes Emergency Generator installed
 6/11

Peralta Homepage, BCC & Merritt website down due to security patch error

8:30AM

10:15AM

8:40AM

10:40AM

  Rebooted Server
 6/5 The PeopleSoft HR and Student Administration modules were unavailable for schedule Maintenance JUNE 5 from 2:00AM June 6 at 8:00AM  Yes Maintenance Completed
 5/17 Server failure  8:45AM 11:20AM   Brought Finance Module up on another server
 5/11 An application server was down.  Emergency reboot of the PeopleSoft Finance module  1:20PM  1:30PM  Yes Rebooted Server
 5/1 Network failed over to secondary firewall due to upgrading of primary firewall software  9:00AM  2:00PM   Moved back to primary firewall

4/26

The PROMT system is temporarily unavailable due to malfunction of primary load balancer 9:45AM 10:00AM   Switch to back up load balancer

4/14

Peralta Homepage, BCC & Merritt website down due to overnight patch error

8:00AM 9:20AM   Rebooted Server

3/31

Windows update caused an network Issue with our DNS 8:00AM 9:20AM   Rebooted Server
 1/14 PeopleSoft HR, BI, & SA Equipment Replacement  5:00PM 8:00PM   Replacement Completed

  1/8

PeopleSoft System  Maintenance 12:00AM   2:00AM   Maintenance Completed

Year 2009

         

  12/18

Cisco Storage Switch & EMC  Storage Processor 4:00AM 4:00AM    

  11/14 & 15

Exchange 2007 Server Maintenance Upgrade 8:00PM 2:00PM   Maintenance Completed

  10/31

Business Intelligence System 3:00PM      

  10/8

PeopleSoft System slow down 10:44AM 1:00PM    

  8/31

Mainframe shut down 4:00PM 4:30PM   Maintenance Completed

  7/29

City of Oakland Electrical Outage 9:00AM     PG&E Replaced Transformer

  6/30

Power Redistribution. Network Services Unavailable 5:00PM 7:00PM   Power Redistributed

  6/3

PG&E Transformer fire. Electrical affected. Network services down 11:45AM 7:00PM   PG&E Replaced Transformer

  5/19

Server Room Electrical Upgrade 9:00AM 9:00PM Yes Electrical Upgrade Completed by Physical Plant

  5/14

  Server Room Electrical Outage – all services affected 12:00PM 3:00PM   Power Restored by Physical Plant

  5/11

Server Room Electrical Outage – all services affected 10:00AM 11:00AM   Power Restored by Physical Plant

  5/8

4 digit dialing and voice mail unavailable for District and Laney College 10:00AM May 9 12:00PM   Advantel Restored service

  4/23

Power outage  in server room. District and Laney College network and telephone services affected 9:00AM 10:00AM   Power Restored by Physical Plant

4/23

Emergency reboot of voice mail equipment due to earlier power outage 2:30PM 2:45PM   Reboot server

4/9

Reset internet connection 12:00PM 12:10PM Yes Reset connection

3/22

Installation of new HVAC and UPS systems in Server Room – Electrical disconnected 6:00AM  3:00PM Yes Install new HVAC and UPS System

3/18

W:Drive maintenance 9:00PM  11:00PM Yes Maintenance Completed

2/14

Server Room Air Conditioning malfunctioned bringing services down 6:00AM  12:00PM Feb 16   Physical Plant repaired HVAC. IT Dept brought services up. One server damaged

1/20

Finance Module unable to access 10:00AM 11:00AM    

1/16

COA Website down for maintenance 3:00PM 5:00PM Yes Patches installed on server
Year 2008          

12/26

EMC Drive Update for PS system and Exchange system 9:00AM 9:00PM Yes Update applied

12/16

AT&T lines down. 7 digit dialing not available 10:00AM 2:30PM   AT&T restored service

12/11

Move PS Servers to faster Cisco switch 12:00PM 12:20PM Yes PS Servers moved to new switch

12/9

College websites moved to faster servers Merritt 10:00AM to 12:30PM BCC 12:30 to 3:00PM Laney 3:00PM to 6:00PM   Yes Websites moved to newer servers

11/10

College websites brought down for emergency maintenance 9:00AM 9:20AM   Servers needed to be rebooted

11/7

Redirect to PROMT system not functioning 4:45PM 5:45PM   Reboot outside DNS

10/31

Peralta website down for maintenance 9:30AM 10:30AM   Maintenance Completed

10/30

Issue accessing PROMT 10:20AM 10:45AM   System rebooted

10/29

Peralta website down for maintenance 7:00PM 8:00PM Yes Server maintenance

10/23

Issues accessing PROMT 10:30AM 12:00PM   Reinstall  PS Application server software on one server

10/22

Sporadic problems accessing PROMT 10:30AM 12:00PM   Application server software bug on one of two applications servers

10/5

PS and Email system maintenance 3:00PM 5:00PM Yes Maintenance Completed

9/26

DNS issue resulting in access issues with services 10:00AM 11:00AM   DNS Server rebooted